OpenStudy (anonymous):

How to write: y=ax^2+bx+c by completing the square and writing it in vertex form ?

OpenStudy (anonymous):

start by setting y = 0 (1) ax^2 + bx + c = 0 c is just a constant, move it to the other side (we're solving for x). (2) ax^2 + bx = -c To complete the square we need the x^2 to be by itself, divide a to get rid of it. (3) x^2 + (b/a)x = -c/a Take half the x term (b/2a) and square it: (b/2a)^2 --->(b^2)/(4a^2) add the last step to both sides of step (3) (4) x^2 + (b/a)x + (b^2)/(4a^2) = -c/a + (b^2)/(4a^2) I'll continue the rest in another post...typing equations sucks.

OpenStudy (anonymous):

oh okay! You can draw it if thast easier![=

OpenStudy (anonymous):

I think I might do that. :)

OpenStudy (anonymous):

ok!!

OpenStudy (anonymous):

Sorry that took long too.

OpenStudy (anonymous):

thxn :DD

OpenStudy (anonymous):

wait but how would you write that in vertex form?

OpenStudy (anonymous):

Sec. I misread your initial question

OpenStudy (anonymous):

oh alrighty ~

OpenStudy (anonymous):

The vertex form of a quadratic is given by y = a(x – h)2 + k, where (h, k) is the vertex. The "a" in the vertex form is the same "a" as in y = ax^2 + bx + c (that is, both a's have exactly the same value). The sign on "a" tells you whether the quadratic opens up or opens down. So...

OpenStudy (anonymous):

Lol. I need to get a new paper out. You better become my fan after this one. :)

OpenStudy (anonymous):

lol ok (x

OpenStudy (anonymous):

I'm going to go ahead an omit the explanation this time. But I will number each step so you can ask me questions on what I did, if you get lost.

OpenStudy (anonymous):

Here you go.

OpenStudy (anonymous):

thnx so very much!

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

&& the axus of s is simply -b/2a=x

OpenStudy (anonymous):

@ajmayberry